OpenStudy (anonymous):

Find the length of the following curve: y=(1/27)(9x^2+6)^3/2 from x=3 to x=6

OpenStudy (anonymous):

arc length?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[L =\int\limits_{a}^{b} ds\]

OpenStudy (anonymous):

So what part is troubling you?

OpenStudy (anonymous):

\[L = \int\limits_{a}^{b} \sqrt{1+[f'(x)]^2} dx\]

OpenStudy (anonymous):

The integration process of the question

OpenStudy (anonymous):

Did you set it up

OpenStudy (anonymous):

\[\huge f(x) = \frac{ 1 }{ 27 }(9x^2+6)^{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

First step, lets find the derivative. Can you do that now please.

OpenStudy (anonymous):

\[L = \int\limits_{3}^{6} \sqrt{1+(x \sqrt{9x^2+6})^{2}} dy\]

OpenStudy (anonymous):

Is this correct so far?

OpenStudy (anonymous):

Yes, that's right. It's better to put it on in later. So with that said, can you square that now, and tell me what you get?

OpenStudy (anonymous):

\[L = \int\limits_{3}^{6} \sqrt{1+(x^2(9x^2 + 6))} \]

OpenStudy (anonymous):

Yes, make sure to add the dx

OpenStudy (anonymous):

Simplify and integrate now.

OpenStudy (anonymous):

ok! Do I simplify the radical now?

OpenStudy (anonymous):

Gotcha

OpenStudy (anonymous):

Inside the radical if you simplify it you'll get \[(3x^2+1)^2\]

OpenStudy (anonymous):

Should be simple from there :P

OpenStudy (anonymous):

Came back to check up on ya, doing good?

OpenStudy (anonymous):

How'd you get (3x^2+1)^2 from that? Just trying to check my simplification

OpenStudy (anonymous):